Per your request I have modified the input coupling to the 2nd Mixer as follows:


         Remove C37

         Replace L15 and L16 with 16 ohm resistors

         Install 0.27 H inductors from pins 2 and 3 of U8 to ground


I then retested the IF passband ripple by sweeping a 1 MHz span centered on 10.7 MHz. The output level was reduced by 3 dB but there was no change in the ripple. Click HERE to view my log.


I think we need to look outside of that coupling area since we’ve pretty much replaced every component. Let’s replace both SAW filters next. What do you think?